OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of three i

jimthompson5910 (jim_thompson5910):

I'm assuming the expression given is \(\Large -3+3\sqrt{3}*i\) If so, then it's the same as last time but now \[\Large a = -3\] \[\Large b = 3\sqrt{3}\]

OpenStudy (anonymous):

yeah, that is the correct expression

OpenStudy (plasmataco):

quick question... what is the square root of i?

OpenStudy (plasmataco):

I know... im a little slow with the noggin but still...

jimthompson5910 (jim_thompson5910):

the i isn't part of the square root

OpenStudy (plasmataco):

i know, just curios.

OpenStudy (anonymous):

-1

OpenStudy (plasmataco):

...

OpenStudy (plasmataco):

i thought that was the square... exponent 2.

OpenStudy (plasmataco):

-1*-1=1 not i.

jimthompson5910 (jim_thompson5910):

\[\Large i = \sqrt{-1}\] \[\Large i^2 = (\sqrt{-1})^2\] \[\Large i^2 = -1\]

OpenStudy (anonymous):

\[\iota ^{2} = -1\]

OpenStudy (plasmataco):

am i just the stupid idiot here or u genuinely dont know or connection problemos?

OpenStudy (plasmataco):

square root. sry like radical.

jimthompson5910 (jim_thompson5910):

sorry @Plasmataco I'm not following

OpenStudy (plasmataco):

i know but like -1 under a radical of 4

jimthompson5910 (jim_thompson5910):

were you able to find r and theta, @lxoser ?

OpenStudy (plasmataco):

...

jimthompson5910 (jim_thompson5910):

oh you mean \[\Large \sqrt{i} = \sqrt[4]{-1}\] @Plasmataco ??

OpenStudy (plasmataco):

\[\sqrt[4]{-1}\]

OpenStudy (plasmataco):

yeah

OpenStudy (anonymous):

im gonna need help finding r and theta @jim_thompson5910

OpenStudy (plasmataco):

sry for interrupting...

OpenStudy (anonymous):

its okay

jimthompson5910 (jim_thompson5910):

\[\Large r = \sqrt{a^2 + b^2}\] \[\Large r = \sqrt{3^2 + (3\sqrt{3})^2}\] ... ... ... \[\Large r = ??\]

OpenStudy (anonymous):

r = 30?

OpenStudy (plasmataco):

dont think so.

OpenStudy (plasmataco):

simplify.

OpenStudy (plasmataco):

it should be somthing a lot lower

jimthompson5910 (jim_thompson5910):

Hint: \[\Large (3\sqrt{3})^2=(3\sqrt{3})*(3\sqrt{3})\] \[\Large (3\sqrt{3})^2=(3*3)*(\sqrt{3}*\sqrt{3})\] \[\Large (3\sqrt{3})^2=(3*3)*\sqrt{3*3}\] \[\Large (3\sqrt{3})^2=9\sqrt{9}\] \[\Large (3\sqrt{3})^2=9*3\] \[\Large (3\sqrt{3})^2=27\]

OpenStudy (anonymous):

r = 6

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (plasmataco):

yup.

OpenStudy (plasmataco):

horay! now my question.

OpenStudy (anonymous):

for theta, i got -60?

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

is the result suppose to be in radians or degrees?

jimthompson5910 (jim_thompson5910):

it depends on what the instructions say

jimthompson5910 (jim_thompson5910):

does it say which mode they want?

OpenStudy (anonymous):

no, but these are the answer choices

jimthompson5910 (jim_thompson5910):

ok so they want radian form

OpenStudy (anonymous):

so would theta be - pi/3

jimthompson5910 (jim_thompson5910):

yes, now because our point is in Q2 (see graph) we add pi radians to our angle to land in the right quadrant 

jimthompson5910 (jim_thompson5910):

- pi/3 is in Q4